A gas whose volume is 450mL is heated from –23 Celsius to 227 Celsius at constant pressure. What is the final volume?
1
Expert's answer
2020-10-15T03:20:52-0400

Applying Charles's Law

"\frac{V_1}{T_1}=\frac{V_2}{T_2}"


"\frac{450}{250}=\frac{V_2}{500}"


"V_2=900 ml"